# Large language models surpass human experts in predicting neuroscience results

### v2

@@ Introduction
Keeping up with the exponentially increasing (?) scientific literature is a superhuman challenge. Potentially disruptive findings go unnoticed in the deluge of articles (?). Processing and integrating the myriad of relevant findings may already surpass humans’ abilities. One path forward is a partnership between human scientists and machines. This partnership could take several forms, including specialist solutions that address specific challenges, such as in protein folding (?), drug discovery (?), and materials science (?). Alternatively, general models of the scientific literature could help guide human scientists’ predictions and study designs. We consider this possibility.
@@
It is an open question whether large language models (LLMs), trained on general text and scientific articles, can predict the outcomes of experiments. If LLMs’ predictions surpassed human experts, the practice of science and the pace of discovery would radically change. We consider this question for neuroscience, which is a large and interdisciplinary field. Prediction in neuroscience should be challenging for human experts for several reasons: (i) there are often many thousands of relevant scientific articles, (ii) an individual study can be noisy, unreliable, and may not replicate, (iii) neuroscience is a multi-level endeavor (?), spanning behavior and molecular mechanisms, (iv) and the analysis methods are diverse and can be complex (?), (v) as are the methods used, which include different brain imaging techniques, lesion studies, gene modification, pharmacological interventions, and so forth.
@@
Can LLMs meet these challenges? In other domains, LLMs have performed impressively. Upon its release, OpenAI’s ChatGPT (?) captured the public’s imagination with its abilities. Most LLMs are based on the transformer architecture (?). These models contain billions and sometimes trillions of weights (?), which are tuned during training in a self-supervised manner to predict the next token, such as the next word in a text passage.
@@
LLMs have displayed remarkable capabilities, including passing professional exams, reasoning (though not without limitations), translation, solving mathematics problems, and even writing computer code (?, ?). By constructing a statistical model during their training to predict the next token, whether that token is a word, pixel, or protein sequence (?), LLMs uncover the underlying patterns or structure of a domain. This generative model captures patterns in the training data, including subtle and imperfect ones. How LLMs learn and generalize to novel situations can be likened to how expert scientists detect patterns in their field after years of reading papers, attending conferences, and analyzing data. From these experiences, human experts build intuitions that enable them to predict future outcomes based on proposed study designs. However, unlike human scientists, LLMs are virtually unconstrained in how much of the scientific literature they can process during training.
@@
How can we formally evaluate the predictive abilities of LLMs in neuroscience? With the rise of LLMs, there has been a surge in evaluation benchmarks, many of which focus on assessing LLMs’ capabilities in scientific domains. Most benchmarks evaluate core knowledge retrieval and reasoning abilities, which are typically backward-looking (Fig. 1). Backward-looking benchmarks include MMLU (?), PubMedQA (?), and MedMCQA (?). These benchmarks are structured in a question-and-answer format, where models must demonstrate extensive world knowledge, retrieve relevant information based on the context of the question, and answer correctly. However, none of these benchmarks are suitable for evaluating the ability of models to predict novel outcomes, which is inherently forward-looking (Fig. 1).
@@
To appreciate how BrainBench qualitatively differs from existing benchmarks, consider a perceived limitation of LLMs, namely their tendency to generate erroneous information, a phenomenon commonly referred to as “hallucination” by LLM researchers. Unlike knowledge graphs that store verified facts, LLMs may not be trustworthy for backward-looking tasks such as summarizing research papers or providing accurate citations (?). However, for forward-looking tasks, such as predicting results from a novel experiment, we view this tendency to mix and integrate information from large and noisy datasets as a virtue. What is a hallucination in a backward-looking task is a generalization or prediction in a forward-looking task (e.g., BrainBench). BrainBench provides a way to quantify this forward-looking ability and compare to human experts. To foreshadow our results, LLMs surpassed human experts on BrainBench by a substantial margin and this margin increased when we provided additional training in neuroscience to an LLM, which we refer to as “BrainGPT”.
@@ Results
Figure 3: Performance of human experts and large language models on BrainBench. (A) LLMs outperformed human experts on BrainBench. Smaller models are on par with larger models. Base versions of models outperformed chat and instruct versions, which were tuned to be conversational with humans. (B) The distribution of test cases across neuroscience subfields roughly mirrors the distribution of articles in the Journal of Neuroscience with Behavior/Cognitive overrepresented. The average performance of $15$ LLMs and human experts is shown. LLMs outperformed human experts in every subfield (see Fig. S.3 for full results). Error bars represent the standard deviation of accuracy around the mean. (C) Majority of the participants were doctoral students, postdoctoral researchers and faculty/academic staff.
@@
On average, LLMs performed better than human experts in every subfield (Fig. 3B), as did each individual LLM (Fig. S.3). Most human experts were doctoral students, postdoctoral researchers, or faculty/academic staff (see Fig. 3C). Please refer to Supplementary materials for more detailed demographic information including years of experience in neuroscience research about the human experts.
@@
When LLMs perform well on a benchmark, one general concern is that the benchmark itself was part of the training set, allowing the LLM to memorize the correct answers. To address this concern, we used a commonly applied measure, zlib-perplexity ratio, for evaluating whether LLMs have memorized passages (?). This ratio gauges the difference between a data-agnostic compression rate of text and data-specific perplexity computed by an LLM (see Methods). Passages that are hard to compress but have low perplexity are indicative of memorization.
We found no indication that BrainBench was memorized by LLMs (Fig. S.5). For comparison, we calculated the zlib-perplexity ratio for a passage that we suspected would be memorized by LLMs, namely the Gettysburg Address. The Gettysburg Address should appear multiple times in an LLM’s training set and indeed it showed signs of memorization (Fig. S.5). Interestingly, for some LLMs, we know exactly what they were trained on (see Table S.2). For these models, the distribution of zlib-perplexity ratios heavily overlapped for items that we knew were in the training set and for items, including BrainBench, that we knew were not in the training set. We believe this overlap indicates that scientific articles, which are unlikely to repeat in training sets, are stored in LLMs as general patterns (akin to a human schema), which support performance on forward-looking tasks (e.g., BrainBench) that require generalization.
@@
To assess whether LLMs’ predictions are calibrated, we examined how well their confidence tracked their accuracy, a crucial characteristic for a trustworthy prediction system. We estimated LLMs’ confidence using the ranked absolute difference in perplexities between two abstracts (Fig. 2; see Methods) and found, like human experts, all LLMs exhibited a positive correlation between accuracy and confidence. When LLMs are confident in their decisions, they are more likely to be correct (Fig. 4).
@@
Pretrained LLMs can provide a foundation for further training in neuroscience with the aim of improving performance, as assessed by BrainBench. We used Low-Rank Adaptation (LoRA) (?) to augment a pretrained LLM, Llama-2-7B-chat, with additional neuroscience knowledge.
LoRA is a parameter-efficient fine-tuning technique that inserts low-rank adapter matrices into LLM transformer blocks (Fig. S.7) and trains only these LoRA weights to update the model’s behavior. In our case, we fine-tuned Llama-2-7B chat using over $1.3$ billion tokens from neuroscience publications spanning $100$ journals between 2002 and 2022 (see Methods), which significantly improved performance.
@@
Figure 5: Fine-tuning a pretrained large language model (LLM) on neuroscience knowledge. Llama-2-7B (chat) was fine-tuned using LoRA on neuroscience articles from 2002-2022 (a total of $1.3$ billion tokens). (A) The fine-tuned model improved by $3\%$ on BrainBench. (B) The fine-tuning process substantially shifted the perplexity distribution of correct responses, indicative of the LLM specializing in neuroscience.
@@ Discussion
We found no indication that LLMs had been exposed to and memorized BrainBench items during their training. Instead, our analyses suggested that LLMs discovered the fundamental patterns that underlie neuroscience studies, which enabled LLMs to predict the outcomes of studies that were novel to them. These conclusions were supported by a widely employed technique (?) to determine text membership within an LLMs’ training set (see Fig. S.5). The Galactica (?) LLMs were particularly illuminating because the contents of their training set are publicly known and did not include BrainBench items. Interestingly, there was no indication of memorization in Galactica for scientific articles that were in its training set, consistent with the notion that LLMs learn broad patterns underlying scientific fields. While passages that frequently repeat in the training set, such as the Gettysburg Address, may be memorized (see Fig. S.5), scientific articles that occur infrequently (most likely once) in the training set appear to support LLM’s forward-looking predictive abilities.
LLM’s impressive forward-looking capabilities suggest a future in which LLMs help scientists make discoveries. To be effective, LLMs need to stay abreast of the rapidly expanding literature. We found that LLMs could efficiently be augmented with neuroscience knowledge using LoRA (?), boosting performance on BrainBench (Fig. 5). LoRA provides a way to create BrainGPT models by re-orienting general-purpose LLMs for use in neuroscience. One can easily imagine a future in which BrainGPT is near continuously updated with new knowledge using LoRA, along with complementary approaches such as Retrieval Augmented Generation (RAG; (?)). RAG could be used to query a database of relevant and up to date scientific articles for the task at hand.
@@
For LLMs to be trustworthy and effective teammates, they need to convey the certainty of their predictions to human scientists. Fortunately, we found that LLMs’ confidence is well calibrated. When LLMs were confident in their predictions, they were more likely to be correct (Fig. 4). A second ingredient for effective teams is being diverse or complementary. LLMs have potential here as well as the items they found difficult did not highly correlate with those human experts found difficult (Fig. S.4). These two ingredients, being well calibrated and complementary, allow systems that combine human and machine judgments to outperform either alone (?).
@@
All our results, including those for calibrated confidence, were only possible because we had access to LLM weights to calculate the perplexity of passages (see Fig. 2). Our approach diverged from the popular approach of prompting models for responses through natural language (i.e., chat). Prompting in natural language may yield less reliable judgments and degrade model competency compared to using model probability scores or training separate classifiers directly from internal representations (?, ?, ?, ?). These observations underscore the importance of working with models that are as open as possible, ideally making both the weights and training set publicly available. Accordingly, we make BrainGPT available on the Huggingface platform https://huggingface.co/BrainGPT.
@@
Beyond serving as a tool for neuroscientists, BrainGPT can help reveal the structure of the field. In particular, we can vary BrainGPT’s training set and observe the effect on BrainBench. For example, what is the effect of including training data from related fields like psychology? In terms of supporting prediction, we can quantify how interrelated fields are. Does it help to weight articles in the training set by their recency, citations, or impact factor? In addition to these training manipulations, we can vary how testing is conducted. For example, would step-by-step thinking via chain-of-thought reasoning (?) benefit BrainGPT? If prediction in neuroscience is akin to a deductive reasoning process, then it should. If instead, as we suspect, prediction in neuroscience is a function of many noisy intertwined signals across subfields, then chain-of-thought reasoning will not help. BrainGPT and BrainBench can help answer these meta-science questions.
@@
We foresee a future in which LLMs serve as forward-looking generative models of the scientific literature. LLMs can be part of larger systems that help researchers determine the best experiment to run next. LLMs’ predictions are informed by a vast scientific literature that no human could read in their lifetime. As LLMs improve, so should their ability to provide accurate predictions. In this contribution, we focused on neuroscience but our aims are broader – we hope to provide a template for any knowledge intensive field. None of the approaches we adopted are neuroscience-specific. We hope to democratize the use of LLMs in science and increase reproducibility by highlighting the use of relatively small models that can be run locally and whose weights are accessible, which contrasts with commercial products. Finally, while LLMs appear poised to supplant humans at prediction, we foresee a role for human experts in providing the accompanying scientific explanations. Prediction is very important, but not everything.

### v3

@@ Introduction
Keeping up with the exponentially increasing  [1] scientific literature is a superhuman challenge. Potentially disruptive findings go unnoticed in the deluge of articles  [2]. Processing and integrating the myriad of relevant findings may already surpass humans’ abilities. One path forward is a partnership between human scientists and machines. This partnership could take several forms, including specialist solutions that address specific challenges, such as in protein folding  [3], drug discovery  [4], and materials science  [5]. Alternatively, general models of the scientific literature could help guide human scientists’ predictions and study designs. We consider this possibility.
@@
It is an open question whether large language models (LLMs), trained on general text and scientific articles, can predict the outcomes of experiments. If LLMs’ predictions surpassed human experts, the practice of science and the pace of discovery would radically change. We consider this question for neuroscience, which is a large and interdisciplinary field. Prediction in neuroscience should be challenging for human experts for several reasons: (i) there are often many thousands of relevant scientific articles, (ii) an individual study can be noisy, unreliable, and may not replicate, (iii) neuroscience is a multi-level endeavor  [6], spanning behavior and molecular mechanisms, (iv) and the analysis methods are diverse and can be complex  [7], (v) as are the methods used, which include different brain imaging techniques, lesion studies, gene modification, pharmacological interventions, and so forth.
@@
Can LLMs meet these challenges? In other domains, LLMs have performed impressively. Upon its release, OpenAI’s ChatGPT  [8] captured the public’s imagination with its abilities. Most LLMs are based on the transformer architecture  [9]. These models contain billions and sometimes trillions of weights  [10], which are tuned during training in a self-supervised manner to predict the next token, such as the next word in a text passage.
@@
LLMs have displayed remarkable capabilities, including passing professional exams, reasoning (though not without limitations), translation, solving mathematics problems, and even writing computer code  [11, 12]. By constructing a statistical model during their training to predict the next token, whether that token is a word, pixel, or protein sequence  [13], LLMs uncover the underlying patterns or structure of a domain. This generative model captures patterns in the training data, including subtle and imperfect ones. How LLMs learn and generalize to novel situations can be likened to how expert scientists detect patterns in their field after years of reading papers, attending conferences, and analyzing data. From these experiences, human experts build intuitions that enable them to predict future outcomes based on proposed study designs. However, unlike human scientists, LLMs are virtually unconstrained in how much of the scientific literature they can process during training.
@@
How can we formally evaluate the predictive abilities of LLMs in neuroscience? With the rise of LLMs, there has been a surge in evaluation benchmarks, many of which focus on assessing LLMs’ capabilities in scientific domains. Most benchmarks evaluate core knowledge retrieval and reasoning abilities, which are typically backward-looking (Fig. 1). Backward-looking benchmarks include MMLU  [14], PubMedQA  [15], and MedMCQA  [16]. These benchmarks are structured in a question-and-answer format, where models must demonstrate extensive world knowledge, retrieve relevant information based on the context of the question, and answer correctly. However, none of these benchmarks are suitable for evaluating the ability of models to predict novel outcomes, which is inherently forward-looking (Fig. 1).
@@
To appreciate how BrainBench qualitatively differs from existing benchmarks, consider a perceived limitation of LLMs, namely their tendency to generate erroneous information, a phenomenon commonly referred to as “hallucination” by LLM researchers. Unlike knowledge graphs that store verified facts, LLMs may not be trustworthy for backward-looking tasks such as summarizing research papers or providing accurate citations  [17]. However, for forward-looking tasks, such as predicting results from a novel experiment, we view this tendency to mix and integrate information from large and noisy datasets as a virtue. What is a hallucination in a backward-looking task is a generalization or prediction in a forward-looking task (e.g., BrainBench). BrainBench provides a way to quantify this forward-looking ability and compare to human experts. To foreshadow our results, LLMs surpassed human experts on BrainBench by a substantial margin and this margin increased when we provided additional training in neuroscience to an LLM, which we refer to as “BrainGPT”.
@@ Results
Figure 3: Performance of human experts and large language models on BrainBench. (A) LLMs outperformed human experts on BrainBench. Smaller models are on par with larger models. Base versions of models outperformed chat and instruct versions, which were tuned to be conversational with humans. (B) The distribution of test cases across neuroscience subfields roughly mirrors the distribution of articles in the Journal of Neuroscience with Behavior/Cognitive overrepresented. The average performance of $15$ LLMs and human experts is shown. LLMs outperformed human experts in every subfield (see Fig. S.5 for full results). Error bars represent the standard deviation of accuracy around the mean. (C) Majority of the participants were doctoral students, postdoctoral researchers and faculty/academic staff. Error bars represent standard error of the mean.
@@
On average, LLMs performed better than human experts in every subfield (Fig. 3B), as did each individual LLM (Fig. S.5). Most human experts were doctoral students, postdoctoral researchers, or faculty/academic staff (see Fig. 3C). Please refer to Supplementary materials for more detailed demographic information including years of experience in neuroscience research about the human experts and distributions of self-reported expertise by subfields (Fig. S.17).
@@
In addition, we analyzed whether LLMs benefited from a general neuroscience context (similar to few-shot prompting) rather than integrating study-relevant information. We tested models using abstracts with sentences randomly swapped from within the same neuroscience subfield. Both original and altered abstracts were used to re-evaluate LLMs’ performance. As shown in Fig. S.4, there was a significant performance decline with coherent versus swapped contexts, indicating that LLMs only partially benefit from accurate, domain-specific, but non-study-relevant context.
@@
When LLMs perform well on a benchmark, one general concern is that the benchmark itself was part of the training set, allowing the LLM to memorize the correct answers. To address this concern, we used a commonly applied measure, zlib-perplexity ratio, for evaluating whether LLMs have memorized passages  [18]. This ratio gauges the difference between a data-agnostic compression rate of text and data-specific perplexity computed by an LLM (see Methods). Passages that are hard to compress but have low perplexity are indicative of memorization.
We found no indication that BrainBench was memorized by LLMs (Fig. S.7). For comparison, we calculated the zlib-perplexity ratio for a passage that we suspected would be memorized by LLMs, namely the Gettysburg Address. The Gettysburg Address should appear multiple times in an LLM’s training set and indeed it showed signs of memorization (Fig. S.7). Interestingly, for some LLMs, we know exactly what they were trained on (see Table S.2). For these models, the distribution of zlib-perplexity ratios heavily overlapped for items that we knew were in the training set and for items, including BrainBench, that we knew were not in the training set. We suspect that the overlap may indicate scientific articles, which are unlikely to repeat in training sets, are stored in LLMs as general patterns, similar to human schemas, supporting performance on tasks requiring generalization (e.g., BrainBench). This hypothesis invites future study.
@@
To assess whether LLMs’ predictions are calibrated, we examined how well their confidence tracked their accuracy, a crucial characteristic for a trustworthy prediction system. We estimated LLMs’ confidence using the ranked absolute difference in perplexities between two abstracts (Fig. 2; see Methods) and found, like human experts, all LLMs exhibited a positive correlation between accuracy and confidence. When LLMs are confident in their decisions, they are more likely to be correct (Fig. 4). In addition, we fitted logistic regressions between model perplexity differences to their correctness as well as human confidences to their correctness on the individual level. We observed significant positive correlations, confirming both models and humans are calibrated (Table S.3).
@@
Pretrained LLMs can provide a foundation for further training in neuroscience with the aim of improving performance, as assessed by BrainBench. We used Low-Rank Adaptation (LoRA)  [19] to augment a pretrained LLM, Mistral-7B-v0.1, with additional neuroscience knowledge.
LoRA is a parameter-efficient fine-tuning technique that inserts low-rank adapter matrices into LLM transformer blocks (Fig. S.18) and trains only these LoRA weights to update the model’s behavior. In our case, we fine-tuned Mistral-7B-v0.1 using over $1.3$ billion tokens from neuroscience publications spanning $100$ journals between 2002 and 2022 (see Methods), which significantly improved performance by 3% on BrainBench (Fig. 5A).
@@
Figure 5: Fine-tuning a pretrained large language model (LLM) on neuroscience knowledge. Mistral-7B-v0.1 was fine-tuned using LoRA on neuroscience articles from 2002-2022 (a total of $1.3$ billion tokens). (A) The fine-tuned model improved by $3\%$ on BrainBench. (B) The fine-tuning process substantially shifted the perplexity distribution of correct responses, indicative of the LLM specializing in neuroscience.
@@ Discussion
We found no indication that LLMs had been exposed to and memorized BrainBench items during their training. Instead, our analyses suggested that LLMs discovered the fundamental patterns that underlie neuroscience studies, which enabled LLMs to predict the outcomes of studies that were novel to them. These conclusions were supported by a widely employed technique  [18] to determine text membership within an LLMs’ training set (see Fig. S.7). The Galactica  [20] LLMs were particularly illuminating because we know which articles were not in the training set vs. ones that might be. Interestingly, there was no indication of memorization in models such as Galactica for scientific articles that were in its training set, consistent with the notion that LLMs learn broad patterns underlying scientific fields. While passages that frequently repeat in the training set, such as the Gettysburg Address, may be memorized (see Fig. S.7), scientific articles that occur infrequently (most likely once) in the training set appear to support LLM’s forward-looking predictive abilities. As a final check, we trained a relatively small LLM from scratch on the published neuroscience literature (excluding preprints and BrainBench items), which eliminated any possible overlap between training data and BrainBench, and found superhuman performance on BrainBench (Fig. S.2).
LLM’s impressive forward-looking capabilities suggest a future in which LLMs help scientists make discoveries. To be effective, LLMs need to stay abreast of the rapidly expanding literature. We found that LLMs could efficiently be augmented with neuroscience knowledge using LoRA  [19], boosting performance on BrainBench (Fig. 5). LoRA provides a way to create BrainGPT models by re-orienting general-purpose LLMs for use in neuroscience. One can easily imagine a future in which BrainGPT is near continuously updated with new knowledge using LoRA, along with complementary approaches such as Retrieval Augmented Generation (RAG;  [17]). RAG could be used to query a database of relevant and up to date scientific articles for the task at hand.
@@
For LLMs to be trustworthy and effective teammates, they need to convey the certainty of their predictions to human scientists. Fortunately, we found that LLMs’ confidence is well calibrated. When LLMs were confident in their predictions, they were more likely to be correct (Fig. 4). A second ingredient for effective teams is being diverse or complementary. LLMs have potential here as well as the items they found difficult did not highly correlate with those human experts found difficult (Fig. S.6). These two ingredients, being well calibrated and complementary, allow systems that combine human and machine judgments to outperform either alone  [21].
@@
All our results, including those for calibrated confidence, were only possible because we had access to LLM weights to calculate the perplexity of passages (see Fig. 2). Our approach diverged from the popular approach of prompting models for responses through natural language (i.e., chat). Prompting in natural language may yield less reliable judgments and degrade model competency compared to using model probability scores or training separate classifiers directly from internal representations  [22, 23, 24, 25]. These observations underscore the importance of working with models that are as open as possible, ideally making both the weights and training set publicly available. Accordingly, we make BrainGPT available on the Huggingface platform https://huggingface.co/BrainGPT.
@@
Beyond serving as a tool for neuroscientists, BrainGPT can help reveal the structure of the field. In particular, we can vary BrainGPT’s training set and observe the effect on BrainBench. For example, what is the effect of including training data from related fields like psychology? In terms of supporting prediction, we can quantify how interrelated fields are. Does it help to weight articles in the training set by their recency, citations, or impact factor? In addition to these training manipulations, we can vary how testing is conducted. For example, would step-by-step thinking via chain-of-thought reasoning  [26] benefit BrainGPT? If prediction in neuroscience is akin to a deductive reasoning process, then it should. If instead, as we suspect, prediction in neuroscience is a function of many noisy intertwined signals across subfields, then chain-of-thought reasoning will not help. BrainGPT and BrainBench can help answer these meta-science questions.
We foresee a future in which LLMs serve as forward-looking generative models of the scientific literature. LLMs can be part of larger systems that assist researchers in determining the best experiment to conduct next. One key step toward achieving this vision is demonstrating that LLMs can identify likely results. For this reason, BrainBench involved a binary-choice between two possible results. LLMs excelled at this task, which brings us closer to systems that are practically useful. In the future, rather than simply selecting the most likely result for a study, LLMs can generate a set of possible results and judge how likely each is. Scientists may interactively use these future systems to guide the design of their experiments.
One risk is that scientists do not pursue studies when their predictions run counter to those of an LLM. In some cases, this might be a sensible course of action, whereas in other cases the LLM might have identified potential gaps or errors in the scientific literature. In the latter situation, conducting the study might result in a significant breakthrough. Conversely, a study result that was predicted with high confidence by an LLM might be viewed as an incremental advance.
LLMs’ predictions are informed by a vast scientific literature that no human could read in their lifetime. As LLMs improve, so should their ability to provide accurate predictions. In this contribution, we focused on neuroscience but our aims are broader – we hope to provide a template for any knowledge intensive field. None of the approaches we adopted are neuroscience-specific. Indeed, the degree of efficacy of our approach may depend on the underlying structure of the domain. For instance, disciplines like mathematics, which rely heavily on logical deduction, might not benefit as much as other scientific fields that involve pattern-based reasoning.
@@
We hope to democratize the use of LLMs in science and increase reproducibility by highlighting the use of relatively small models that can be run locally and whose weights are accessible, which contrasts with commercial products. Finally, while LLMs appear poised to supplant humans at prediction, we foresee a role for human experts in providing the accompanying scientific explanations. Prediction is very important, but not everything.
Data availability
Human participant data, and intermediate data generated via simulations and analyses are publicly available at https://github.com/braingpt-lovelab/BrainBench. Model weights and training data are available at https://huggingface.co/BrainGPT.
Code availability
All computer code associated with this work including model training, evaluation, data processing and analyses are publicly available at https://github.com/braingpt-lovelab/BrainBench.